# Pre-Training and Personalized Fine-Tuning via Over-the-Air Federated Meta-Learning: Convergence-Generalization Trade-Offs

### v5

@@ I Introduction
For modern artificial intelligence (AI) applications such as large language models (LLMs), the training paradigm has shifted to pre-training followed by fine-tuning [1]. Pre-training is currently done centrally using large data repositories obtained, often at a high cost, by large corporations [2, 1, 3]; while fine-tuning is typically much cheaper, and it can be used to personalize models to individual requirements [4, 5]. For example, a foundation LLM can be fine-tuned to serve as a personal assistant or as a surrogate CEO based on data tailored to the individual use cases[5]. As large dataset repositories are becoming a scarce commodity, it has become imperative to use distributed data in a privacy-minded way for pre-training and, possibly, fine-tuning [6, 7]. Furthermore, efforts to democratize access to AI models are moving in the direction of federated pre-training implementations, which leverage decentralized computing resources [8].
@@
A general framework in which pre-training and fine-tuning can be formalized is meta-learning. In meta-learning, data from different tasks are used to pre-train a model with the aim of ensuring that the pre-trained model can be efficiently fine-tuned based on limited data for a new task [9, 10]. When applied to a multi-agent, and federated setting, as illustrated in Fig. 1, agents collaborate via communication to a central server for pre-training of a shared model defined by hyperparameters $\theta$, from which each agent $i$ can extract a fine-tuned model $\mbox{{$\varphi$}}_{i}$ that is tailored to the individual data sets at each agent $i$. As in federated learning (FL), agents do not directly exchange data but only model information, thus limiting the leakage of private information [8].
@@
Conventional FL, which aims at finding a single shared model for all agents, is known to suffer from the heterogeneity of the distributions across agents. Heterogeneity has been addressed via methods such as decomposition [11], masking [12], and clustering [13]. Personalized FL (pFL) alleviates the problem of data heterogeneity by optimizing personalized AI models that are adjusted to the local tasks of interest for each user [14].
@@
Meta-learning-based pFL (meta-pFL) moves beyond basic personalization by targeting generalization to new agents and tasks [15, 16, 17]. In line with the goal of the pre-training/fine-tuning workflow, meta-pFL does not merely aim for enhancing the performance of existing agents and tasks. Rather, it optimizes shared hyperparameters $\theta$ that can be efficiently personalized via fine-tuning to new agents and tasks that are not observed during federated pre-training.
@@
This paper studies the generalization performance of meta-pFL for a wireless setting in which the agents participating in the pre-training phase, i.e., meta-learning, are connected via a shared wireless channel to the server [18, 19]. Wireless meta-pFL systems are suitable for large-scale intelligent Internet of Things (IoT) systems, in which IoT devices collect data for collaborative model training. However, the performance of such systems may suffer from limited access to wireless resources. A popular class of methods has explored the use of over-the-air computing [20], an approach known as over-the-air FL (AirFL) [20, 18, 19], to overcome communication bottlenecks on the shared radio channel for conventional FL.
@@
Channel impairments may not necessarily be deleterious for generalization. In fact, the addition of noise during pre-training can have a regularizing effect [21], and it may even implement approximate forms of Bayesian learning [10] [22]. Accordingly, references such as [23, 24, 25] have shown improved generalization performance as a result of transmissions of model information over noisy channels in AirFL. The downside of a noisy training process is the convergence time, which may be severely affected by noisy updates [26, 27, 28]. This paper studies the trade-off between generalization to new agents and tasks, on the one hand, and convergence, on the other hand, for an over-the-air implementation of meta-pFL.
@@ I-A Related Works
AirFL protocols for conventional FL have been widely studied. Among some notable algorithmic contributions, the authors in [19, 29] proposed and optimized truncated-channel-inversion power control strategies, while reference [30] proposed to enhance spectral efficiency via sparsification and linear compression mechanisms. Convergence was studied in  [26, 27] for phase compensation-based or channel inversion-based AirFL schemes targeting convex loss functions, showing a linear convergence rate as in the centralized gradient descent (GD). These studies reveal the impact of channel fading, channel noise, and data heterogeneity on the convergence performance of AirFL. Other related works on FL include [31, 28], which derived and optimized transmission powers to minimize a convergence error upper bound of AirFL.
Several pFL protocols have been proposed that apply in the presence of ideal channels. For instance, some methods leverage knowledge distillation (KD) to obtain personalized models by transferring knowledge of a powerful model to local models [32]; while others divide the trainable model parameters into personalized and shared parts [11]. The meta-pFL protocol studied in [15] builds on the model agnostic meta-learning paradigm (MAML) [33, 34], which applies a two-level gradient descent approach to learn how to quickly adapt to new tasks. Reference [15] analyzed the convergence of meta-pFL in the presence of ideal communication.
Limited work has been done on the design of pFL protocols for wireless systems. Solutions include [35, 36], cluster-based pFL [37, 38] and partial-model-based pFL [39], which are based on digital transmission. The analysis in these papers provides insights into convergence. AirFL-based schemes for pFL were considered in [40, 41, 42], focusing on convergence analysis and optimal resource allocation.
For meta-pFL, the generalization error is a key performance metric, as it quantifies the capacity of a shared model to be efficiently fine-tuned to new tasks. Generalization analyses for conventional AirFL can be found in [25], which shows that heavy-tail noise deteriorates the convergence rate while improving the generalization capacity. References [43, 44] have studied the generalization error for MAML under the assumption of ideal channels. To the best of our knowledge, no analysis of generalization exists for meta-pFL under the assumption of ideal channels, let alone noisy channels.
@@ I-B Contributions
    We introduce a MAML-based over-the-air meta-pFL protocol, which adapts the approach in [15] for use of shared wireless channels. The proposed protocol, termed Air-meta-pFL, leverages sparsification and linear compression [30] along with a long-term memory mechanism to compensate for the error caused by the gradient sparsification [45].
@@ II Preliminaries and System Model
In this section, we first review pFL and meta-pFL [15], which operate under the assumption of ideal and noiseless communications. Then, we describe the considered wireless channel model that accounts for fading and noise. Throughout this work, we study the setting in Fig. 1, in which a set $n]\triangleq\{1,\ldots,n\}$ of devices communicate with an edge server over a multiple access (MAC) fading channel.
@@ II-A Preliminaries
where $\hat{\nabla}f_{i}(\mbox{{$\theta$}}_{i}^{(t,q)})=\frac{1}{|\mathcal{B}_{i}^{(t)}|}\sum_{Z\in\mathcal{B}_{i}^{(t)}}\nabla\ell(\mbox{{$\theta$}}_{i}^{(t,q)};Z)$ is an estimate of the gradient of the local test loss (1) obtained from a mini-batch $\mathcal{B}_{i}^{(t)}\subseteq\mathcal{D}_{i}^{\rm(tr)}$ of data points; and the vector $\hat{\nabla}f_{i}(\mbox{{$\theta$}}_{i}^{(t,q)}-\alpha\hat{\nabla}f_{i}(\mbox{{$\theta$}}_{i}^{(t,q)}))$ and Hessian matrix $\hat{\nabla}^{2}f_{i}(\mbox{{$\theta$}}_{i}^{(t,q)})$ are similarly obtained from distinct mini-batches of $\mathcal{D}_{i}^{\rm(va)}$. Note that the estimate $\hat{\nabla}F_{i}(\mbox{{$\theta$}})$ is not unbiased, i.e., $\mathbb{E}[\hat{\nabla}F_{i}(\mbox{{$\theta$}})]\neq\nabla F_{i}(\mbox{{$\theta$}})$ [15].
@@ II-B Communication Model
where $\mbox{{$n$}}^{(t)}$ is the additive white Gaussian noise (AWGN) vector with zero mean and variance $\sigma_{n}^{2}$. We also assume that the interference from non-AirFL IoT devices is negligible or can be eliminated via signal processing methods [46].
As in [30], we assume ideal and noiseless communication in the downlink, given the less constrained resources available for downlink communication at the edge server.
@@ III-A Air-meta-pFL
As described in the previous section, in meta-pFL at each communication round $t$, each active device $i\in\mathcal{I}^{(t)}$ must communicate its local update $\mbox{{$\Delta$}}_{i}^{(t)}$ in (8) to the server. To this end, as in [30, 26, 47, 28], we adopt an analog communication strategy based on sparsification, linear compression, channel phase compensation, and power scaling.
@@
First, a $k$-contraction operator is applied to sparsify vector $\mbox{{$\Delta$}}_{i}^{(t)}$. The $k$-contraction operator $\text{Comp}_{k}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ preserves only $k$ entries of its input vector, while setting all other entries to zeros, e.g., via a top-$k$ selection or via a random selection strategy [45]. The resulting sparsified update is given by
@@
where a memory vector $\mbox{{$m$}}_{i}^{(t)}\in\mathbb{R}^{d}$ is used to keep track of the accumulated errors as [45]
@@ IV-A Assumptions
Assumption 1 ($k$-Contraction [45])
@@
Assumption 4 (Bounded Data Heterogeneity [15])
@@
Note that Assumption 5 is trivially satisfied for a scheme that does not implement sparsification and linear compression by setting the variance of the estimation error $v^{(t)}$ to be equal to the channel noise power, i.e., $v^{(t)}=\sigma^{2}$ for all $t$. In the more general case, this assumption is satisfied by applying a linear minimum mean squared error (LMMSE) estimator. It is also approximately met by Bayesian sparse recovery algorithms [48], for which one can evaluate the estimation error $v^{(t)}$ by leveraging state evolution [49, 48].
@@ V-A Meta-Generalization Error
As in the generalization analysis of MAML [44], we study the generalization performance of Air-meta-pFL in terms of the local test loss experienced by a new device that did not participate in the federated learning process. Recalling that each device is characterized by a data distribution $\mu\sim\tau$, let us write the corresponding meta-test loss (2) for the new device directly as a function of the data-generation distribution $\mu$ as the expectation
@@ V-B Assumptions
Assumption 7 (Sub-Gaussian Per-task Training Loss [43])
@@
Assumption 7 is automatically satisfied with $\sigma^{2}=(b-a)^{2}/4$ if the loss function $\ell(\cdot;\cdot)$ satisfies the inequality $a\leq\ell(\cdot;\cdot)\leq b$. In practice, one can always satisfy the sub-Gaussian assumption by clipping the original loss or by applying a composition with a squashing function such as sigmoid function [50].
@@ V-C Generalization Analysis
To start, we review the following lemma derived in [43], which bounds the meta-generalization error in (30) for any protocol described by a conditioned distribution $\mathrm{P}_{\boldsymbol{\theta}|\mathcal{D}_{1:n}}$ via the mutual information $I(\mbox{{$\theta$}};\mathcal{D}_{1:n})$ between the pre-trained vector $\theta$ and the training data set $\mathcal{D}_{1:n}$. Note that conventional techniques to analyze generalization performance, such as VC dimension or Rademacher complexity, yield bounds that are determined by the complexity of the model class rather than training algorithms. By comparison, the mutual information is algorithm-dependent [44], yielding stronger design insights.
Lemma V.1 (Theorem 1, [43])
@@ VI Numerical Results
Meta-learning rate $\eta^{(t)}=\eta=0.4$ Estimator $\mathcal{E}(\cdot)$ OAMP with $20$ iterations [48]
@@ VI-A Experimental Settings
We adopt the Omniglot data set, which contains $1643$ characters, considered as classes, with each class having $p=20$ instances drawn by different persons, amounting to $32860$ instances in total [33, 44, 51]. We employ $n=9$ devices for pFL training and $n_{\text{test}}=3$ additional devices for evaluation of the generalization performance. All devices are assigned non-$i.i.d.$ local data sets.
The experimental setting is illustrated in Fig. 3. Specifically, our experiments adopt the $N$-way $K$-shot classification protocol as follows [33]. Note that first-order MAML achieves comparable test accuracy as MAML for $N$-way $K$-shot tasks [33, Table 1] (see Fig. 6), which provides evidence that the proposed method with first-order approximation is potentially applicable to large models. The task of each training device $i\in[n]$ is to distinguish among $N$ different classes, selected out of a set of $0\leq m_{c}\leq 1643$ classes specific to each device, given $K$ instances of each of the $N$ classes. Performance is then assessed by the test accuracy for classification of instances on newly $N$ selected classes within the device-specific $m_{c}$ classes. For each training device $i\in[n]$, $m_{c}$ different characters (classes) are drawn uniformly without replacement from Omniglot. The data set $\mathcal{D}_{i}$ includes $p=20$ instances for each of the $m_{c}$ classes. The local data sets for the $n_{\text{test}}$ test devices are generated following the same procedure. Accordingly, the level of data heterogeneity across devices decreases with $m_{c}$. In our experiment, we set $N=5$ and $K=8$. At each communication round $t$, each active device $i\in\mathcal{I}^{(t)}$ randomly chooses a mini-batch $m_{B}=32$ NN-way $K$-shot tasks from data set $\mathcal{D}_{i}$ for inner update and another $m_{B}=32$ tasks for outer update. The tasks are obtained by selecting $N$ classes uniformly at random from the $m_{c}$ available classes. The meta-test loss $\bar{F}_{\tau}(\mbox{{$\theta$}})$ and test accuracy are averaged over $1000$ different $N$-way $K$-shot tasks of each test device, and over $n_{\text{test}}$ test devices.

### v6

@@ I Introduction
For modern artificial intelligence (AI) applications such as large language models (LLMs), the training paradigm has shifted to pre-training followed by fine-tuning [touvron2023llama]. Pre-training is currently done centrally using large data repositories obtained, often at a high cost, by large corporations [achiam2023gpt, touvron2023llama, team2023gemini]; while fine-tuning is typically much cheaper, and it can be used to personalize models to individual requirements [hu2021lora, wei2021finetuned]. For example, a foundation LLM can be fine-tuned to serve as a personal assistant or as a surrogate CEO based on data tailored to the individual use cases[wei2021finetuned]. As large dataset repositories are becoming a scarce commodity, it has become imperative to use distributed data in a privacy-minded way for pre-training and, possibly, fine-tuning [sun2025federated, ni2025federated]. Furthermore, efforts to democratize access to AI models are moving in the direction of federated pre-training implementations, which leverage decentralized computing resources [mcmahan2017communication].
@@
A general framework in which pre-training and fine-tuning can be formalized is meta-learning. In meta-learning, data from different tasks are used to pre-train a model with the aim of ensuring that the pre-trained model can be efficiently fine-tuned based on limited data for a new task [chen2023learning, simeone2022machine]. When applied to a multi-agent, and federated setting, as illustrated in Fig. 1, agents collaborate via communication to a central server for pre-training of a shared model defined by hyperparameters $\theta$, from which each agent $i$ can extract a fine-tuned model $\mbox{{$\varphi$}}_{i}$ that is tailored to the individual data sets at each agent $i$. As in federated learning (FL), agents do not directly exchange data but only model information, thus limiting the leakage of private information [mcmahan2017communication].
@@
Conventional FL, which aims at finding a single shared model for all agents, is known to suffer from the heterogeneity of the distributions across agents. Heterogeneity has been addressed via methods such as decomposition [arivazhagan2019federated], masking [setayesh2022perfedmask], and clustering [mansour2020three]. Personalized FL (pFL) alleviates the problem of data heterogeneity by optimizing personalized AI models that are adjusted to the local tasks of interest for each user [Tan2022towards].
@@
Meta-learning-based pFL (meta-pFL) moves beyond basic personalization by targeting generalization to new agents and tasks [fallah2020personalized, wang2023memory, khodak2019adaptive]. In line with the goal of the pre-training/fine-tuning workflow, meta-pFL does not merely aim for enhancing the performance of existing agents and tasks. Rather, it optimizes shared hyperparameters $\theta$ that can be efficiently personalized via fine-tuning to new agents and tasks that are not observed during federated pre-training.
@@
This paper studies the generalization performance of meta-pFL for a wireless setting in which the agents participating in the pre-training phase, i.e., meta-learning, are connected via a shared wireless channel to the server [yang2020federated, zhu2019broadband]. Wireless meta-pFL systems are suitable for large-scale intelligent Internet of Things (IoT) systems, in which IoT devices collect data for collaborative model training. However, the performance of such systems may suffer from limited access to wireless resources. A popular class of methods has explored the use of over-the-air computing [nazer2007computation], an approach known as over-the-air FL (AirFL) [nazer2007computation, yang2020federated, zhu2019broadband], to overcome communication bottlenecks on the shared radio channel for conventional FL.
@@
Channel impairments may not necessarily be deleterious for generalization. In fact, the addition of noise during pre-training can have a regularizing effect [russo2019much], and it may even implement approximate forms of Bayesian learning [simeone2022machine] [welling2011bayesian]. Accordingly, references such as [liu20privacy, liu2022wireless, yang2021revisiting] have shown improved generalization performance as a result of transmissions of model information over noisy channels in AirFL. The downside of a noisy training process is the convergence time, which may be severely affected by noisy updates [sery2020analog, sery2021over, wen2023convergence]. This paper studies the trade-off between generalization to new agents and tasks, on the one hand, and convergence, on the other hand, for an over-the-air implementation of meta-pFL.
@@ I-A Related Works
AirFL protocols for conventional FL have been widely studied. Among some notable algorithmic contributions, the authors in [zhu2019broadband, WCNC2023fullversion] proposed and optimized truncated-channel-inversion power control strategies, while reference [amiri2020federated] proposed to enhance spectral efficiency via sparsification and linear compression mechanisms. Convergence was studied in  [sery2020analog, sery2021over] for phase compensation-based or channel inversion-based AirFL schemes targeting convex loss functions, showing a linear convergence rate as in the centralized gradient descent (GD). These studies reveal the impact of channel fading, channel noise, and data heterogeneity on the convergence performance of AirFL. Other related works on FL include [cao2021optimized, wen2023convergence], which derived and optimized transmission powers to minimize a convergence error upper bound of AirFL.
Several pFL protocols have been proposed that apply in the presence of ideal channels. For instance, some methods leverage knowledge distillation (KD) to obtain personalized models by transferring knowledge of a powerful model to local models [li2019fedmd]; while others divide the trainable model parameters into personalized and shared parts [arivazhagan2019federated]. The meta-pFL protocol studied in [fallah2020personalized] builds on the model agnostic meta-learning paradigm (MAML) [finn2017model, rajeswaran2019meta], which applies a two-level gradient descent approach to learn how to quickly adapt to new tasks. Reference [fallah2020personalized] analyzed the convergence of meta-pFL in the presence of ideal communication.
Limited work has been done on the design of pFL protocols for wireless systems. Solutions include [yue2022efficient, you2023automated], cluster-based pFL [li2022energy, zhao2023ensemble] and partial-model-based pFL [chen2023knowledge], which are based on digital transmission. The analysis in these papers provides insights into convergence. AirFL-based schemes for pFL were considered in [chen2023personalizing, mortaheb2022personalized, sami2022over], focusing on convergence analysis and optimal resource allocation.
For meta-pFL, the generalization error is a key performance metric, as it quantifies the capacity of a shared model to be efficiently fine-tuned to new tasks. Generalization analyses for conventional AirFL can be found in [yang2021revisiting], which shows that heavy-tail noise deteriorates the convergence rate while improving the generalization capacity. References [jose2021information, chen2021generalization] have studied the generalization error for MAML under the assumption of ideal channels. To the best of our knowledge, no analysis of generalization exists for meta-pFL under the assumption of ideal channels, let alone noisy channels.
@@ I-B Contributions
    We introduce a MAML-based over-the-air meta-pFL protocol, which adapts the approach in [fallah2020personalized] for use of shared wireless channels. The proposed protocol, termed Air-meta-pFL, leverages sparsification and linear compression [amiri2020federated] along with a long-term memory mechanism to compensate for the error caused by the gradient sparsification [stich2018sparsified].
@@ II Preliminaries and System Model
In this section, we first review pFL and meta-pFL [fallah2020personalized], which operate under the assumption of ideal and noiseless communications. Then, we describe the considered wireless channel model that accounts for fading and noise. Throughout this work, we study the setting in Fig. 1, in which a set $n]\triangleq\{1,\ldots,n\}$ of devices communicate with an edge server over a multiple access (MAC) fading channel.
@@ II-A Preliminaries
where $\hat{\nabla}f_{i}(\mbox{{$\theta$}}_{i}^{(t,q)})=\frac{1}{|\mathcal{B}_{i}^{(t)}|}\sum_{Z\in\mathcal{B}_{i}^{(t)}}\nabla\ell(\mbox{{$\theta$}}_{i}^{(t,q)};Z)$ is an estimate of the gradient of the local test loss (1) obtained from a mini-batch $\mathcal{B}_{i}^{(t)}\subseteq\mathcal{D}_{i}^{\rm(tr)}$ of data points; and the vector $\hat{\nabla}f_{i}(\mbox{{$\theta$}}_{i}^{(t,q)}-\alpha\hat{\nabla}f_{i}(\mbox{{$\theta$}}_{i}^{(t,q)}))$ and Hessian matrix $\hat{\nabla}^{2}f_{i}(\mbox{{$\theta$}}_{i}^{(t,q)})$ are similarly obtained from distinct mini-batches of $\mathcal{D}_{i}^{\rm(va)}$. Note that the estimate $\hat{\nabla}F_{i}(\mbox{{$\theta$}})$ is not unbiased, i.e., $\mathbb{E}[\hat{\nabla}F_{i}(\mbox{{$\theta$}})]\neq\nabla F_{i}(\mbox{{$\theta$}})$ [fallah2020personalized].
@@ II-B Communication Model
where $\mbox{{$n$}}^{(t)}$ is the additive white Gaussian noise (AWGN) vector with zero mean and variance $\sigma_{n}^{2}$. We also assume that the interference from non-AirFL IoT devices is negligible or can be eliminated via signal processing methods [jha2023analog].
As in [amiri2020federated], we assume ideal and noiseless communication in the downlink, given the less constrained resources available for downlink communication at the edge server.
@@ III-A Air-meta-pFL
As described in the previous section, in meta-pFL at each communication round $t$, each active device $i\in\mathcal{I}^{(t)}$ must communicate its local update $\mbox{{$\Delta$}}_{i}^{(t)}$ in (8) to the server. To this end, as in [amiri2020federated, sery2020analog, xing2021federated, wen2023convergence], we adopt an analog communication strategy based on sparsification, linear compression, channel phase compensation, and power scaling.
@@
First, a $k$-contraction operator is applied to sparsify vector $\mbox{{$\Delta$}}_{i}^{(t)}$. The $k$-contraction operator $\text{Comp}_{k}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ preserves only $k$ entries of its input vector, while setting all other entries to zeros, e.g., via a top-$k$ selection or via a random selection strategy [stich2018sparsified]. The resulting sparsified update is given by
@@
where a memory vector $\mbox{{$m$}}_{i}^{(t)}\in\mathbb{R}^{d}$ is used to keep track of the accumulated errors as [stich2018sparsified]
@@ IV-A Assumptions
Assumption 1 ($k$-Contraction [stich2018sparsified])
@@
Assumption 4 (Bounded Data Heterogeneity [fallah2020personalized])
@@
Note that Assumption 5 is trivially satisfied for a scheme that does not implement sparsification and linear compression by setting the variance of the estimation error $v^{(t)}$ to be equal to the channel noise power, i.e., $v^{(t)}=\sigma^{2}$ for all $t$. In the more general case, this assumption is satisfied by applying a linear minimum mean squared error (LMMSE) estimator. It is also approximately met by Bayesian sparse recovery algorithms [ma2017orthogonal], for which one can evaluate the estimation error $v^{(t)}$ by leveraging state evolution [donoho2009message, ma2017orthogonal].
@@ V-A Meta-Generalization Error
As in the generalization analysis of MAML [chen2021generalization], we study the generalization performance of Air-meta-pFL in terms of the local test loss experienced by a new device that did not participate in the federated learning process. Recalling that each device is characterized by a data distribution $\mu\sim\tau$, let us write the corresponding meta-test loss (2) for the new device directly as a function of the data-generation distribution $\mu$ as the expectation
@@ V-B Assumptions
Assumption 7 (Sub-Gaussian Per-task Training Loss [jose2021information])
@@
Assumption 7 is automatically satisfied with $\sigma^{2}=(b-a)^{2}/4$ if the loss function $\ell(\cdot;\cdot)$ satisfies the inequality $a\leq\ell(\cdot;\cdot)\leq b$. In practice, one can always satisfy the sub-Gaussian assumption by clipping the original loss or by applying a composition with a squashing function such as sigmoid function [rivasplata2012subgaussian].
@@ V-C Generalization Analysis
To start, we review the following lemma derived in [jose2021information], which bounds the meta-generalization error in (30) for any protocol described by a conditioned distribution $\mathrm{P}_{\boldsymbol{\theta}|\mathcal{D}_{1:n}}$ via the mutual information $I(\mbox{{$\theta$}};\mathcal{D}_{1:n})$ between the pre-trained vector $\theta$ and the training data set $\mathcal{D}_{1:n}$. Note that conventional techniques to analyze generalization performance, such as VC dimension or Rademacher complexity, yield bounds that are determined by the complexity of the model class rather than training algorithms. By comparison, the mutual information is algorithm-dependent [chen2021generalization], yielding stronger design insights.
Lemma V.1 (Theorem 1, [jose2021information])
@@ VI Numerical Results
Meta-learning rate $\eta^{(t)}=\eta=0.4$ Estimator $\mathcal{E}(\cdot)$ OAMP with $20$ iterations [ma2017orthogonal]
@@ VI-A Experimental Settings
We adopt the Omniglot data set, which contains $1643$ characters, considered as classes, with each class having $p=20$ instances drawn by different persons, amounting to $32860$ instances in total [finn2017model, chen2021generalization, lake2011one]. We employ $n=9$ devices for pFL training and $n_{\text{test}}=3$ additional devices for evaluation of the generalization performance. All devices are assigned non-$i.i.d.$ local data sets.
The experimental setting is illustrated in Fig. 3. Specifically, our experiments adopt the $N$-way $K$-shot classification protocol as follows [finn2017model]. Note that first-order MAML achieves comparable test accuracy as MAML for $N$-way $K$-shot tasks [finn2017model, Table 1] (see Fig. 6), which provides evidence that the proposed method with first-order approximation is potentially applicable to large models. The task of each training device $i\in[n]$ is to distinguish among $N$ different classes, selected out of a set of $0\leq m_{c}\leq 1643$ classes specific to each device, given $K$ instances of each of the $N$ classes. Performance is then assessed by the test accuracy for classification of instances on newly $N$ selected classes within the device-specific $m_{c}$ classes. For each training device $i\in[n]$, $m_{c}$ different characters (classes) are drawn uniformly without replacement from Omniglot. The data set $\mathcal{D}_{i}$ includes $p=20$ instances for each of the $m_{c}$ classes. The local data sets for the $n_{\text{test}}$ test devices are generated following the same procedure. Accordingly, the level of data heterogeneity across devices decreases with $m_{c}$. In our experiment, we set $N=5$ and $K=8$. At each communication round $t$, each active device $i\in\mathcal{I}^{(t)}$ randomly chooses a mini-batch $m_{B}=32$ NN-way $K$-shot tasks from data set $\mathcal{D}_{i}$ for inner update and another $m_{B}=32$ tasks for outer update. The tasks are obtained by selecting $N$ classes uniformly at random from the $m_{c}$ available classes. The meta-test loss $\bar{F}_{\tau}(\mbox{{$\theta$}})$ and test accuracy are averaged over $1000$ different $N$-way $K$-shot tasks of each test device, and over $n_{\text{test}}$ test devices.